# Systematic bias in LISA ringdown analysis due to waveform inaccuracy

### v1

@@ Appendix A Reliability of the linear signal regime
We see that when $N<3$, although relative waveform errors are at the percent level, the maximum dephasing is close to 1, suggesting that the linear signal approximation is not valid. Recall that one of the features that we observed in Fig. 3 was the oscillatory behavior of $\Delta^{(\rm Sys)}\theta^{i}$ with increasing $N$ for $N<3$. Such an oscillatory behavior is tied to the breakdown of the linear signal approximation. We can see this through the oscillatory $\exp(i\Delta\Psi)$ dependence in $\Delta^{(\rm Sys)}\theta^{i}$ (when one does not further linearize $\Delta^{(\rm Sys)}\theta^{i}$ in $\Delta\Psi$). As $N$ increases to $63\leq N<$ 6, we observe that $\Delta\Psi_{\max}$ drops to $\mathcal{O}(10^{-1})$, and we enter the regime where the linear signal approximation is valid. Crucially, for $N\geq 6$, $\Delta\Psi_{\max}$ is $\mathcal{O}(10^{-2})$, ensuring that the linear signal approximation is indeed valid in estimating the minimum number of modes, and that our estimate of $N_{\min}=7$ is robust. While we demonstrated this validity for a particular system in Fig. 5, the arguments apply across parameter space. The agreement between our estimate of $N_{\min}$ based on the (explicit linear signal approximation) criteria in Eq. (23) and the (implicit linear signal approximation) criteria in Eq. (25) further strengthens the robustness and self-consistency obtained within the linear signal approximation.
@@ Appendix B Comparison with NR waveforms
$h_{\ell m}^{Fit}=\sum_{\ell\textquoteright=-\ell}^{\ell\textquoteright=\ell}\mu_{\ell\ell\textquoteright m\textquoteright n\textquoteright}h_{\ell\textquoteright m\textquoteright n\textquoteright}\delta_{mm\textquoteright},$ (30)
[…]
where $\mu_{\ell\ell\textquoteright m^{\prime}n^{\prime}}$ are the spherical-spheroidal mixing coefficients Berti and Klein (2014); London and Fauchon-Jones (2019). Note that in this expression, the linear perturbation framework is respected and only the modes that fulfill $m\textquoteright=m$ are considered. To give some context, modes with $m\textquoteright\neq m$ (known as ‘recoil’ modes) are present in several NR waveforms due to extrapolation at null infinity. If, instead, one extracts the numerical waveforms with the Cauchy-characteristic extraction technique in the superrest (BMS) frame Moxon et al. (2021); Mitman et al. (2022); Magaña Zertuche et al. (2022), these recoil modes vanish.
@@ Appendix C The impact of spins, mass ratio and angles on $N_{\text{min}}$
[…]
In Fig. 8, we show the dependence of $N_{\text{min}}$ on the spin (upper panel) and mass ratio (lower panel) of the progenitors, for a system with primary mass $m_{1}=5\times 10^{6}M_{\odot}$ and luminosity distance $dL=10\,\text{Gpc}$. The angles are fixed to $\theta=\psi=\iota=\pi/3$, $\phi=0$, and we consider equal spins. The discontinuous behavior that can be observed in both panels reflects the evolution of the modes ordering. In particular, the sudden jumps correspond to the exchange of two or more modes in the SNR ranking. We do not observe a clear trend of $N_{\text{min}}$ in any of the two cases. Overall, the impact of spin and mass ratio is mainly on the ordering of the modes, and only changes $N_{\text{min}}$ of ±$\pm 1$ mode, a smaller fluctuation with respect to the variability introduced by the orientation angles and shown in Fig. 7. For very loud sources requiring a large $N_{\text{min}}$, variations due to spin and mass ratio become larger but remain mostly subdominant compared to those from sky location and inclination.
@@ Appendix D A mode-dependent exponential tapering in the frequency domain
Thus, our phenomenological tapering mimics a faster than exponential fall-off (due to $1/f^{2}$ and $1/f^{3}$ for the $+$ and $\times$ polarizations respectively). This approach is conceptually similar to the frequency-domain tapering of IMR waveforms Husa et al. (2016); Khan et al. (2016); Mehta et al. (2023).
The comparison of these exponentially tapered functions and the standard combinations of Lorentzian functions is shown in Fig. 9 for the $220$ mode and a BH mass of $10^{6}\,M_{\odot}$.
Our tapering can be motivated by constructing a time-domain Planck window McKechan et al. (2010) for each mode. With a Planck window $W_{\rm Planck}(t)$, the high frequency fall-off is at most exponential, which can be sketched out in the following way. The Fourier transform of a $C^{\infty}$ window function $W(t)$ with compact support, upon integration by parts $p$ times, can be expressed as $W(f)=W^{(p)}(f)/(i2\pi f)^{p}$, where $W^{(p)}(f)$ is the Fourier transform of the $p$-th derivative of $W(t)$. Since every derivative exists due to $W(t)$ being $C^{\infty}$ and $W(t)$ has compact support, we have that $|W(f)|\leq A/f^{p}$, suggesting that $W(f)$ decays faster than every polynomial in frequency. In other words, $|W(f)|$ can at most have an exponential fall-off. This implies that the $C^{\infty}$ Planck window will satisfy $|W_{\rm Planck}(f)|\leq A\exp(-Bf)$, where $B\sim\Delta t$ with $\Delta t$ being the characteristic transition time of the window . With our phenomenological tapering, for each mode, the corresponding mode-dependent Planck window would then have an exponential fall-off $B_{\ell mn}\sim 1/\omega_{\ell mn}$. However, by tuning the mode-dependent transition time $\Delta t_{\ell mn}$, in practice, one can get a faster exponential fall-off with a mode-dependent Planck window.
@@
which is also equivalent to Taylor expanding $g(t)$. Given that $W(f)=W^{(p)}(f)/(i2\pi f)^{p}$, to leading order for the frequency fall-off, we simply have that $|W(f)|\sim 1/f^{p+1}$. For the commonly used Heaviside window for ringdown analysis, this means that the spectral leakage fall-off is only $1/f$. Instead, we use a frequency domain ringdown model obtained from mirroring Flanagan and Hughes (1998); Berti et al. (2006b) at $t=t_{0}$. Although mirroring is not identical to a window function multiplying the damped sinusoid, due to the mirrored waveform being $C^{0}$ at $t=t_{0}$, it mimics the use of a $C^{0}$ window, and thus leads to a $1/f^{2}$ fall-off (cf. Eq. (II.1)). Note, however, that $\Lambda^{\times}_{\ell mn}$ scales as $1/f^{3}$ instead of $1/f^{2}$ as is the case for $\Lambda^{+}_{\ell mn}$. This is due to the fact that $\Lambda^{\times}_{\ell mn}$ is an anti-symmetric combination which makes the leading $1/f^{2}$ term cancel out, leaving the next-to-leading $1/f^{3}$ scaling.
@@
For a source with remnant mass and spin of $2.66\times 10^{7}\,M_{\odot}$ and $0.60$ respectively, and angles $\theta=3\pi/4$, $\phi=\psi=0$, $\iota=\pi/3$ and $\varphi=\pi$, we apply the mismatch criterion given by Eq. (25) with the two methods (time and frequency domain). We show the result in Fig. 10. The mismatch $1-{\cal M}$ is shown by solid lines, while its threshold (right-hand side of Eq. (25)) is shown by dashed lines. The comparison is presented for different models, namely: the time domain (TD) mismatch in red; the frequency domain mismatch with a common single cutoff in purple; and the frequency domain mismatch with optimal tapering in blue. With the cutoff mismatch, two modes are sufficient for a correct estimation of the parameters, while seven modes are needed with the optimal taper. For the time-domain mismatch, four modes are sufficient to fulfill this requirement. We can therefore conclude that with the single cutoff, we can conservatively estimate the minimum number of modes needed for unbiased parameter estimation with damped-sinusoids waveforms.

### v2

@@ Appendix A Reliability of the linear signal regime
We see that when $N<3$, although relative waveform errors are at the percent level, the maximum dephasing is close to 1, suggesting that the linear signal approximation is not valid. Recall that one of the features that we observed in Fig. 3 was the oscillatory behavior of $\Delta^{(\rm Sys)}\theta^{i}$ with increasing $N$ for $N<3$. Such an oscillatory behavior is tied to the breakdown of the linear signal approximation. We can see this through the oscillatory $\exp(i\Delta\Psi)$ dependence in $\Delta^{(\rm Sys)}\theta^{i}$ (when one does not further linearize $\Delta^{(\rm Sys)}\theta^{i}$ in $\Delta\Psi$). As $N$ increases to $3\leq N<6$, we observe that $\Delta\Psi_{\max}$ drops to $\mathcal{O}(10^{-1})$, and we enter the regime where the linear signal approximation is valid. Crucially, for $N\geq 6$, $\Delta\Psi_{\max}$ is $\mathcal{O}(10^{-2})$, ensuring that the linear signal approximation is indeed valid in estimating the minimum number of modes, and that our estimate of $N_{\min}=7$ is robust. While we demonstrated this validity for a particular system in Fig. 5, the arguments apply across parameter space. The agreement between our estimate of $N_{\min}$ based on the (explicit linear signal approximation) criteria in Eq. (23) and the (implicit linear signal approximation) criteria in Eq. (25) further strengthens the robustness and self-consistency obtained within the linear signal approximation.
@@ Appendix B Comparison with NR waveforms
$h_{\ell m}^{Fit}=\sum_{\ell’=-\ell}^{\ell’=\ell}\mu_{\ell\ell’m’n’}h_{\ell’m’n’}\delta_{mm’},$ (30)
Figure 7: Dependence of $N_{\text{min}}$ on the cosine of the angles $\theta$ and $\iota$, for three values of the mass ratio of the progenitors $q$ and $a_{1}=a_{2}=0$. The primary BH mass and the luminosity distance have been fixed respectively to $5\times 10^{6}M_{\odot}$ and $5\,\text{Mpc}$. $N_{\text{min}}$ is represented by the color code. Observe that the color pattern is symmetric with respect to $\cos\theta=0$, while the corresponding symmetry about $\cos\iota=0$ is slightly broken. Observe also that some patterns present in the right panel disappear in the middle one, to appear again in the mass-symmetric case (left panels). This is the consequence of the competing effects at play when the progenitor mass ratio changes, namely (de)excitation of odd-$m$ modes against changes in the global SNR.
where $\mu_{\ell\ell’m^{\prime}n^{\prime}}$ are the spherical-spheroidal mixing coefficients Berti and Klein (2014); London and Fauchon-Jones (2019). Note that in this expression, the linear perturbation framework is respected and only the modes that fulfill m=mm’=m are considered. To give some context, modes with $m’\neq m$ (known as ‘recoil’ modes) are present in several NR waveforms due to extrapolation at null infinity. If, instead, one extracts the numerical waveforms with the Cauchy-characteristic extraction technique in the superrest (BMS) frame Moxon et al. (2021); Mitman et al. (2022); Magaña Zertuche et al. (2022), these recoil modes vanish.
@@ Appendix C The impact of spins, mass ratio and angles on $N_{\text{min}}$
In Fig. 7, we show the dependence of $N_{\text{min}}$ on the angles $\iota$ and $\theta$, fixing the progenitors spins to $a_{1}=a_{2}=0$ and q=0.1, 0.5, 1q=0.1,\,0.5,\,1. First, we notice that the patterns are completely symmetric around $\cos\theta=0$, whereas there is a slight symmetry breaking in the two hemispheres defined by $\cos\iota=0$. This is indeed what we expect, as two configurations with opposite sky localization with respect to the horizon are completely equivalent, while the corresponding symmetry in the BH frame is broken by the direction of the spin. Furthermore, it can be observed that the case with $q=0.5$ represents an optimal configuration, where more modes are needed on average, compared to the other two cases. This happens because, as we decrease the mass ratio starting from $q=1$, the QNM spectrum becomes richer, as odd-$m$ modes are switched on. On the other hand, as we go to smaller mass ratios while keeping the primary mass fixed, the size of the region in which $N_{\text{min}}<5$ increases again. This happens because we are decreasing the total SNR of the event, although many modes are excited. For $q=0.1$ there are still, however, marginal regions in which $N_{\text{min}}>6$.
Figure 8: Dependence of $N_{\text{min}}$ on progenitor spins and mass ratio. The primary BH mass and the luminosity distance have been fixed respectively to $5\times 10^{6}M_{\odot}$ and $10\,\text{Gpc}$, while the angles have been fixed to $\theta=\psi=\iota=\pi/3$, $\phi=0$. The jumps correspond to crossing of two or more modes in the single-mode SNR ordering.
In Fig. 8, we show the dependence of $N_{\text{min}}$ on the spin (upper panel) and mass ratio (lower panel) of the progenitors, for a system with primary mass $m_{1}=5\times 10^{6}M_{\odot}$ and luminosity distance $d_{L}=10\,\text{Gpc}$. The angles are fixed to $\theta=\psi=\iota=\pi/3$, $\phi=0$, and we consider equal spins. The discontinuous behavior that can be observed in both panels reflects the evolution of the modes ordering. In particular, the sudden jumps correspond to the exchange of two or more modes in the SNR ranking. We do not observe a clear trend of $N_{\text{min}}$ in any of the two cases. Overall, the impact of spin and mass ratio is mainly on the ordering of the modes, and only changes $N_{\text{min}}$ of $\pm 1$ mode, a smaller fluctuation with respect to the variability introduced by the orientation angles and shown in Fig. 7. For very loud sources requiring a large $N_{\text{min}}$, variations due to spin and mass ratio become larger but remain mostly subdominant compared to those from sky location and inclination.
@@ Appendix D A mode-dependent exponential tapering in the frequency domain
Thus, our phenomenological tapering mimics a faster than exponential fall-off (due to $1/f^{2}$ and $1/f^{3}$ for the ++ and $\times$ polarizations respectively). This approach is conceptually similar to the frequency-domain tapering of IMR waveforms Husa et al. (2016); Khan et al. (2016); Mehta et al. (2023). The comparison of these exponentially tapered functions and the standard combinations of Lorentzian functions is shown in Fig. 9 for the $220$ mode and a BH mass of $10^{6}\,M_{\odot}$.
Our tapering can be motivated by constructing a time-domain Planck window McKechan et al. (2010) for each mode. With a Planck window $W_{\rm Planck}(t)$, the high frequency fall-off is at most exponential, which can be sketched out in the following way. The Fourier transform of a $C^{\infty}$ window function $W(t)$ with compact support, upon integration by parts $p$ times, can be expressed as $W(f)=W^{(p)}(f)/(i2\pi f)^{p}$, where $W^{(p)}(f)$ is the Fourier transform of the $p$-th derivative of $W(t)$. Since every derivative exists due to $W(t)$ being $C^{\infty}$ and $W(t)$ has compact support, we have that $|W(f)|\leq A/f^{p}$ (with $A$ being an overall integration factor), suggesting that $W(f)$ decays faster than every polynomial in frequency. In other words, $|W(f)|$ can at most have an exponential fall-off. This implies that the $C^{\infty}$ Planck window will satisfy $|W_{\rm Planck}(f)|\leq A\exp(-Bf)$, where $B\sim\Delta t$, with $\Delta t$ being the characteristic transition time of the window. With our phenomenological tapering, for each mode, the corresponding mode-dependent Planck window would then have an exponential fall-off $B_{\ell mn}\sim 1/\omega_{\ell mn}$. However, by tuning the mode-dependent transition time $\Delta t_{\ell mn}$, in practice, one can get a faster exponential fall-off with a mode-dependent Planck window.
@@
which is also equivalent to Taylor expanding $g(t)$. Given that $W(f)=W^{(p)}(f)/(i2\pi f)^{p}$, to leading order for the frequency fall-off, we simply have that $|W(f)|\sim 1/f^{p+1}$. For the commonly used Heaviside window [^7] Note that although the Heaviside window doesn’t vanish for $t\rightarrow\infty$, the result still applies because $h(t)W(t)$ is compact throughout the domain, with $h(t)$ being the GW strain. for ringdown analysis, this means that the spectral leakage fall-off is only $1/f$. Instead, we use a frequency domain ringdown model obtained from mirroring Flanagan and Hughes (1998); Berti et al. (2006b) at $t=t_{0}$. Although mirroring is not identical to a window function multiplying the damped sinusoid, due to the mirrored waveform being $C^{0}$ at $t=t_{0}$, it mimics the use of a $C^{0}$ window, and thus leads to a $1/f^{2}$ fall-off (cf. Eq. (II.1)). Note, however, that $\Lambda^{\times}_{\ell mn}$ scales as $1/f^{3}$ instead of $1/f^{2}$ as is the case for $\Lambda^{+}_{\ell mn}$. This is due to the fact that $\Lambda^{\times}_{\ell mn}$ is an anti-symmetric combination which makes the leading $1/f^{2}$ term cancel out, leaving the next-to-leading $1/f^{3}$ scaling.
@@
[…]